An arrow moving at 10 m/s is stopped by a target. If the arrow has a mass of 250 g and penetrates 4 cm into the target, what is the force exerted by the arrow on the target?


1
Expert's answer
2022-05-17T16:04:49-0400

The energy-work theorem says

"\frac{mv^2}{2}=Fd"

"F=\frac{mv^2}{2d}=\frac{0.250*10^2}{2*0.04}=313\:\rm N"